OpenStudy (anonymous):

what is the sum of a 7-term geometric series if the first term is -11the last term is -171,875 and the common ratioi -5

OpenStudy (campbell_st):

ok... so the sum of a geometric series is \[S_{n} = \frac{a(1 - r^n)}{1 - r}\] you know n = 7, a = -11 and r = -5 just substitute and calculate the answer

OpenStudy (anonymous):

which is a again

OpenStudy (campbell_st):

sorry I don't have my calculator...

OpenStudy (anonymous):

no to substitute

OpenStudy (campbell_st):

a is the 1st term... -11

OpenStudy (anonymous):

ILL LET U KNOW WAT I GOT

OpenStudy (anonymous):

-13020

OpenStudy (campbell_st):

umm that seems low \[S_{7} = \frac{-11 \times (1 - (-5)^7)}{1 - (-5)}\] I have to go... good luck with it